APPRENTICE

Name (24)  @uchicago.edu
Emily Bargar  ebargar
Piotr S. Behr  pbehr
Alexander Belyi  abelyi
Oleg Chanyshev  olegch
Chingwei Chyr  chyr
Corrin Clarkson  clarkson
Arpit Gupta  arpit
Thomas J. Hansberger  thomash
Eric Janofsky  ebjan
Robert A. Kreek  rakreek
Alice Mark  amark
Thomas Monson  currence
Jacqueline Mussach  jpeg
Hae Young Noh  hynoh
Yana Peysakhovich  yana
Jeffrey Picel  jpicel
Nikhil Raghuram  nraghura
Rebecca Resnick  rebeccar
James Salvatore  sal13
Cyrus Shirzadi  cyrus
Maxwell Shron  mshron
Lisa Wang  lywang
David Yu  dryu
Wayne (Weizhen) Zhong  waynez

SPECIAL SIT-IN

Name (2)  @uchicago.edu
Emily Norton  brntnrtn
Sara Rezvi  arsinoe

YSP

Name (22)  Program  @uchicago.edu
Sam Altschul  11-12  salt
Semere G. Baraki  7-8  sgb
Gabriel Bender  11-12  gbender
Matthew Bruce  7-8  mbruce
Christopher Chariker 9-10  chariker
Sarah Constantin  7-8  sconstan@Princeton.EDU
Stephen David Flood  7-8  sflood
Nicolas Ford  9-10  nford
Daniel Gardner  11-12  dbgardner
Gabriel Gaster  9-10  eulerian
Joseph Gunther  9-10  jgunther
Shawn Henry  11-12  shawnh
Phillip Kalikman  11-12  plk
Zhiwei Calvin Lin  9-10  linzw
Jeff Meyer  7-8  jsmeyer
Jonathan Nieder  9-10  jrnieder
Samuel Raskin  9-10  sraskin
Edward T. Ruggeri  7-8  ruggeri
Jeff Schonert  11-12  jeffrey1
Abigail Sheldon  7-8  sheldona
Weiwei Tan  9-10  cora7
Brian Taylor  7-8  bgtaylor

SESAME

Name (26)  @uchicago.edu
Stephen Balady  sbalady
Padraic James Bartlett  padraic
Chad Casarotto  chadmbol
Trevor Davis  trevorld
Daniel Arthur Dickinson  dand
Charles Doss  cdoss
Andy Eisenberg  aeisen
Matthew Gonwa  gonwa
Sayer Herin  sayerh
Gregory J. Herschlag  Gtzu
Drew Hudek  dhudec
Daniel Immerman  immerman
Milan Kidd  mkidd
Andrew Kiersz  akiersz
Asa Levi  asa
Joshua C. McNeur  jmcneur
Nilsen Miller  nilsen1
Matthew O’Meara  momeara
Robert Prag  rprag
Sarah Raskin  sarahr
Katrina Redelsheimer  kredelsheimer
Andrew Steiner  steinera
John Sternberg  jds3
Michael Thvedt  cynicalk
Muhammead Waliji  mhwaliji
Dan Weiner  dbw